OpenStudy (anonymous):

Let f and g be the functions given by f(x)=e^x and g(x) = ln x. Find the area of the region enclosed by the graphs of f and g between x = 1/2 and x=1.

myininaya (myininaya):

\[\int\limits\limits_{.5}^{1} (upper function-lower function) dx\]

OpenStudy (anonymous):

on that interval \[\ln(x)\] is negative and \[e^x\] is positive, so if you want the area between them you can integrate \[-\int_{\frac{1}{2}}^1\ln(x)dx+\int_{\frac{1}{2}}^1e^xdx\]

myininaya (myininaya):

yep upper function is the e^x and the lower function is ln(x)

OpenStudy (anonymous):

or what myininaya says. it amounts to the same thing

OpenStudy (anonymous):

the only hard one is the first one. \[\int_.5^1e^xdx=e-\sqrt{e}\]

OpenStudy (anonymous):

'anti derivative" of log is \[x\ln(x)-x\] so that one is not so hard either

OpenStudy (anonymous):

there are 2 parts to this question...can u help me with te hsecond part too?

OpenStudy (anonymous):

so is the final asnwer xln(x) - x + (e - sqrt of x)